Fix gross profit (loss) in requirement 5 only

DONT USE THIS WAY

Gross profit recognized = revenue -actual cost

2018 3,000,000-2,400,000 = 600,000

2019 3,019,400-3,800,000 = -780,600

2020 -100,000 actual loss - profit recognized in 2018 - loss recognized in 2019

= -100,000 - 600,000 - (-780,600)

=80600

2020 gross profit = 80600

I POSTED THIS QUESTION TWICE BEFORE AND THEY USED THIS METHOD AND IT IS NOT CORRECT

IT IS WRONG I DONT WANT TO SEE IT AGAIN

I DONT WANT TO SEE -300,000 ON 2019

I DONT WANT TO SEE -100,000 ON 2020

ALL WRONG AS SHOWN

In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows:
FIX

2018 2019 2020
Cost incurred during the year $ 2,400,000 $ 3,600,000 $ 2,200,000
Estimated costs to complete as of year-end 5,600,000 2,000,000 0
Billings during the year 2,000,000 4,000,000 4,000,000
Cash collections during the year 1,800,000 3,600,000 4,600,000


Westgate recognizes revenue over time according to percentage of completion.

Required:
1. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years.
2-a. In the journal below, complete the necessary journal entries for the year 2018 (credit "Various accounts" for construction costs incurred).
2-b. In the journal below, complete the necessary journal entries for the year 2019 (credit "Various accounts" for construction costs incurred).
2-c. In the journal below, complete the necessary journal entries for the year 2020 (credit "Various accounts" for construction costs incurred).
3. Complete the information required below to prepare a partial balance sheet for 2018 and 2019 showing any items related to the contract.
4. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information.

2018 2019 2020
Cost incurred during the year $ 2,400,000 $ 3,800,000 $ 3,200,000
Estimated costs to complete as of year-end 5,600,000 3,100,000 0


5. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information.

2018 2019 2020
Cost incurred during the year $ 2,400,000 $ 3,800,000 $ 3,900,000
Estimated costs to complete as of year-end 5,600,000 4,100,000 0


Answer is complete but not entirely correct. Complete this question by entering your answers in the tabs below Req 3 Req 5 Req 2A Req 2B Req 2C Req 1 Req 4 Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years. (Do not round intermediate calculations. Loss amounts should be indicated with a minus sign.) 2018 2019 2020 Revenue 3,000,0004,500,000 2,500,000 Gross profit 600,000 900,000 300,000 OSs Req 2A> Req 1
In the journal below, complete the necessary journal entries for the year 2018 (credit "Various accounts" for construction costs incurred). (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.) Date General Journal Debit Credit 0 2019 Construction in progress 2,400,000 2400,000 Various accounts 2019 2,000,000 Acco unts receivable Billings on construction contract 2,000,000 2019 Cash 1,800,000 unts receivable 1,800,000 4 2019 Cost of construction 2,400,000 Construction in progress 600,000 3,000,000 Revenue from long-term contracts
In the journal below, complete the necessary journal entries for the year 2019 (credit "Various accounts" for construction costs incurred). (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.) Date General Journal Debit Credit 2019 Construction in progress 3,600,000 3,600,000 Various accounts 4,000,000 2019 Accounts receivable Billings on construction contract 4,000,000 3,600,000 2019 Cash 3,600,000 Accounts receivable 3,600,000 4 2019 Cost of construction Construction in progress 900,000 Revenue from long-term contracts 4,500,000
In the journal below, complete the necessary journal entries for the year 2020 (credit "Various accounts" for construction costs incurred). (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.) Debit Date General Journal Credit 2020 Construction in progress 2,200,000 2.200.000 Various accounts 2020 Acco unts receivable 4,000,000 4,000,000 Billings on construction contract 2020 Cash 4,600,000 unts receivable 4,600,000 >| 2,200,000 4 2020 Cost of construction Construction in progress 300,000 Revenue from long-term contracts 2,500,000
Req 3 Req 1 Req 2A Req 2C Req 4 Req 2B Req 5 Complete the information required below to prepare a partial balance sheet for 2018 and 2019 showing any items related to the contract. (Do not round intermediate calculations.) Balance Sheet (Partial) 2018 2019 Current assets $ 600,000 Accounts receivable $ 200,000 S 3,000,000 (2,000,000) $7,500,000 (6,000,000) Construction in progress Less: Billings Costs and profit in excess of billings 1,000,000 1,500,000 Current liabilities Req 4 K Req 2C
Answer is complete but not entirely correct Complete this question by entering your answers in the tabs below. Req 2A Req 1 Req 2B Req 2C Req 3 Req 4 Req 5 Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount. Loss amounts should be indicated with a minus sign.) 2020 2018 2019 3,000,000 3,667,000 3,333,000 Revenue Gross proft loss) 600,000 (133,333) 133,333 K Req 3 Req 5 >
Answer is complete but not entirely correct. Complete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req 4 Req 2C Req 3 Req 5 Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount. Loss amounts should be indicated with a minus sign.) 2018 2019 2020 3,000,000 3,019,400 3,980,600 600,000 (300,000) 100,000 Revenue Gross profit oss
Answer #1

5)

2018 2019 2020
estimated cost to complete [A] 5600000 4100000 0
cost incurred till date[B] 2400000 2400000+3800000= 6200000 6200000+3900000=10100000
Total cost to complete [A+B] 8000000 10300000 10100000

**Cost incurred till date = Cost incurred till last period+ cost incurred during the current year

2018 2019 2020
% of completion [cost incurred till date /total cost to complete] 2400000/8000000 6200000/10300000 10100000/10100000
30% 60.19417% 100%
2018 2019 2020
Revenue to recognized {(Contract price *%of completion)-revenue recognized till date} {(10,000,000*30%)-0}=3000000

{(10000000*60.19417%)-3000000}

{6019400-3000000)

3019400

{(10,000,000*100%)-6019400}

{10,000,000-6019400}

3980600

Gross profit (Revenue recognized -cost incurred during the year) 3000000-2400000= 600000 -900000 200000

***since in 2019 ,the total cost to complete exceeds contract price ,any loss on contract should be immediately recognized as per accounting principle .

Loss on contract = 10000000-10300000= -300000

Gross profit /(loss) to recognize in year 2019 = profit/(loss) on contract - any profit recognized in earlier years

                            = -300000- 600000

                           = -900000

2020;

Profit/(loss) on contract = 10000000-10100000= -100000

Gross profit /(loss) to recognized in 2020 = -100000-(-300000)

                                                  = -100000 +300000

                                                     = 200000